Mr. D.B. Carpenter Flat 8, Broughton House 28/30 Bullar Road Bitterne Park Southampton, S018 1GS

January 5, 1996

Vanpac International Ltd, Unit 16, Central Trading Estate, Marine Parade, Southampton SO14 5JP

Dear Vanpac,

I enclose the acceptance form for your estimate, together with the insurance proposal, the US customs declaration, the Power of Attorney, the supplementary customs declaration, and photocopies of the relevant pages of my passport.

I also enclose a cheque for £653.58 as agreed.

My initial day-time contact number in Syracuse, USA will be 315-443-1722. I will let you know if this changes.

Yours sincerely,